# A galactic outflow traced by its extended Mg II emission out to a $\sim 30$ kpc radius in the Hubble Ultra Deep Field with MUSE

### v1

@@ 3.2 Integrated spectrum
In Fig. 2 we present a spectrum of the galaxy extracted by summing the datacube over a circular aperture of 1″ radius. The spectrum displays a blue stellar continuum with prominent Balmer absorption and strong nebular emission of [O II], [Ne III], the Balmer series, [O III], and other well-known lines. The spectrum shows Mg II to be mainly in absorption, but with significant redshifted P-Cygni emission humps as revealed by the inset panel in Fig. 2. Furthermore, Mg I $\lambda$2852 is detected in absorption.
@@
The emission line spectrum of #884 is that of a normal star-forming galaxy of moderate excitation level. The line widths also reveal no evidence for an AGN, nor do we find any AGN-characteristic emission line such as [Ne V] $\lambda$3426 in the spectrum. This is in agreement with the marginal detection ($5\sigma$, in the soft band only) as an X-ray source in the 7 Ms exposure of the Chandra observatory (Luo et al., 2017). The X-ray luminosity of $2\times 10^{41}$ erg s-1 is an order of magnitude below the limit usually adopted as the criterion for AGN, in line with the soft X-ray spectrum. Given the observed star formation rate, it is also consistent with the expected X-ray emission arising from stellar processes alone (Mineo et al., 2014). While these diagnostics do not exclude the possiblity of a weak AGN in the center of the galaxy, all indicators suggest that any such source would have to be energetically subdominant.
Figure 3: Mg II pseudo-narrow band image extracted from the MUSE datacube by summing over the spectral range of the doublet and filtering with a Gaussian of 1″ FWHM. Underlying in blue scale is the HST F775W image, with other galaxies labeled as in Fig. 1. The black solid contours are spaced logarithmically at 0.25 dex interval, starting at $\log_{10}(s)$ of $-18.5$, where $s$ is surface brightness in erg s-1 cm-2 arcsec-2. The lowest contour corresponds to a S/N of $\sim$1.5 in the filtered image. The dotted contours indicate negative measured values of $s$. The central hole in the Mg II surface brightness distribution is a result of the continuum subtraction, since #884 is a net absorber in its central spectrum. Similarly, the apparently negative region at the background galaxy BG1 is caused by the strong Mg II absorption in the continuum of BG1. The colored circumferences indicate the radii of the annular apertures used to extract radially resolved spectra in Sec. 4.2.
@@ 4.1 Mg II narrow band image
Similarly to the central depression, the Mg II emission in Fig. 3 also seems to be avoiding the galaxy BG1, and possibly also FG1. While for the former the explanation is again strong Mg II absorption in the underlying continuum of BG1, this is not possible for FG1 as the object is located in the foreground. However, any intrinsic stellar absorption feature in the FG1 spectrum that happens to be located within the Mg II narrow band range will have a similar effect. Since BG1 is sufficiently far away from #884 in terms of projected distance ($\sim 30$ kpc), we keep BG1 and its immediate surroundings outside the region where we model the Mg II emission (see Sec. 7).
At the upper edge of the image we see the bright foreground galaxy FG2 and the physical neighbour NB1. Most of the apparent signal in these spaxels is due to continuum-subtraction residuals from FG2 that are not restricted to the Mg II wavelength range. Similarly to BG1, this galaxy is also beyond the modeled region, to avoid our results being influenced by contaminant emission.
On the other hand, we find that NB1 is not surrounded by any large-scale extended Mg II emission (its apparent halo in Fig. 3 is an artifact generated by the residual of the continuum subtraction of FG2). The spectrum of NB1 shows only weak Mg II emission, with a 2-to-1 ratio between the two Mg II emission lines, suggesting a nebular origin. However, it is certainly possible that #884 and NB1 are in gravitational interaction, and that this interaction could influence the morphology of the Mg II halo through tidal forces, in the sense that it shows its largest extent along the direction towards NB1 (see, e.g., Leclercq et al., 2022).
Figure 4: Spatially resolved Mg II spectra of galaxy #884, extracted as the sum of spaxels in circular concentric ring-like apertures (circular aperture in the case of the central region). Each panel corresponds to a different annulus as indicated in the label, except for the top-left panel, which corresponds to the spectrum of the full modeled FoV. In the frame of reference of the central galaxy, the apertures represent physical sizes of $<3$ kpc, $3-6$ kpc, $6-9$ kpc, $9-15$ kpc, and $15-26$ kpc. The error bars represent the ±$\pm 1\sigma$ uncertainty of the spectra. The nearly horizontal dotted lines indicate the approximate continuum level. The vertical thin dashed lines mark the observed wavelengths of the Mg II doublet at the systemic redshift of the galaxy.
@@ 4.2 Radially resolved spectra
In Fig. 4 we show spectra of the Mg II doublet, extracted in five disjoint concentric circular annular apertures, defined by the radii $<0\aas@@fstack{\prime\prime}4$, $0\aas@@fstack{\prime\prime}4-0\aas@@fstack{\prime\prime}8$, $0\aas@@fstack{\prime\prime}8-1\aas@@fstack{\prime\prime}2$, and $1\aas@@fstack{\prime\prime}2-2\aas@@fstack{\prime\prime}0$, or $<3$ kpc, $3-6$ kpc, $6-9$ kpc, $9-15$ kpc, and $15-26$ kpc, in the frame of reference of the central galaxy, as well as for the full modeled FoV. These apertures are shown in Fig. 3, following the same color scheme as in Fig. 4. The central spectrum (magenta) is dominated by absorption with only very weak P-Cygni emission humps. The emission then gets progressively stronger compared to the absorption when moving outwards. At $1\aas@@fstack{\prime\prime}2<r<2\aas@@fstack{\prime\prime}0$ and beyond the absorption is essentially undetected.
@@ 5 Model
The main goal of this paper is to design a semianalytic model, with only few parameters to avoid overfitting, that can reproduce key features observed of the Mg II halo of #884, in terms of the spatial distribution and line profiles of the Mg II emission and absorption, and that can be fitted to spatially resolved IFU observations. In the following subsections, we describe in detail the different components and other relevant aspects considered for our model, in order to achieve this goal. Figure 5 provides a flow chart that summarizes our modeling approach. Further in the paper, we describe the fit of our model to the data, aiming to unveil some of the key physics at play in the circumgalactic gas.
Figure 5: Flow chart that summarizes our modeling and fitting scheme. Gray rounded boxes correspond to inputs for the model, blue rectangles correspond to intermediate outputs, and purple hexagons correspond to processes. The final output is indicated with a red ellipse. For a given set of parameters, a model cube is constructed on a spaxel-by-spaxel basis. The iteration over different individual spaxels is indicated with the yellow area. We use an MCMC routine to find the best-fitting parameters that match our data cube. The complete MCMC iteration over the parameter space is enclosed within the green (including the yellow) area. The inputs and intermediate outputs surrounded by a thick box are static, that is, they do not change over any of the iterations performed. In the top-left part of some boxes, we indicate the section where the concept inside the box is described. See Secs. 5 and 6 for a comprehensive description of our modeling and fitting approach.
@@ 5.1 Basic model
In this section, we describe the outflow model introduced by Scuderi et al. (1992) in the context of stellar winds and further developed by Scarlata & Panagia (2015), Carr et al. (2018), and Carr et al. (2021) to model the scattered resonant emission in expanding envelopes around galaxies. We also explain how we adapt this modeling scheme to integral field spectroscopic data. The basic model consists of a spherically symmetric source of radius $R_{0}$ isotropically emitting continuum radiation, surrounded by a spherical envelope of gas (wind) expanding from radius $R_{0}$ up to a terminal radius $R_{\rm max}$. The velocity at which the spherical envelope expands ($v$) increases with radius ($r$), according to a power law of exponent $\gamma$, such that:
@@
where $v_{0}$ corresponds to the initial velocity of the wind (i.e., at $R_{0}$), and $v_{\mathrm{max}}$ is the terminal velocity of the wind at $R_{\mathrm{max}}$. A wind velocity that increases with radius is consistent with results from simulations (see, e.g., Cooper et al., 2008), and it is also consistent with the down-the-barrel observations reported by Martin & Bouché (2009), although different, more complex velocity profiles are possible (see, e.g., Murray et al., 2011). The density profile corresponding to Eq. 1, assuming mass conservation through the outflow, is:
$n(r)=n_{0}\,\biggl{(}\frac{R_{0}}{r}\biggr{)}^{\gamma+2}$ (2)
@@
where $n_{0}$ is the gas density at $r=R_{0}$. Provided the velocity gradient in Eq. 1 is large, the photons produced by the central source will interact with the outflowing material only at the specific radius where the absorbing ions are at resonance (due to their Doppler shift), an assumption commonly known as the “Sobolev” approximation. However, due to projection effects, the line-of-sight velocity $v_{\rm obs}$ measured by an observer for a given spherical shell at radius $r$ will differ from its real velocity, and will depend on the projected distance ($\epsilon$) of the line of sight to the central source (a.k.a., impact parameter), going from $v(r)$ at the center ($\epsilon=0$), to zero at $\epsilon=r$ (see also Scarlata & Panagia, 2015; Carr et al., 2018, for a more detailed explanation). On the other hand, it can be shown that the surfaces of constant $v_{\rm obs}$ do not cross each other, and hence, for any given line of sight, there is a bijective relation between frequency (or velocity) space and spatial location within the expanding envelope (because any sightline intersects different isovelocity surfaces only once). Figure 6 shows the shape of the (observed) isovelocity curves for $\gamma$ = 0.5, 1.0, and 1.5.
@@
Figure 6: Isovelocity curves in the spherical outflow model. The curves are calculated following Eq. 3 for three different values of $\gamma$, setting $R_{0}$ and $v_{0}$ to 1. The observer sits far away from the object towards the positive z-direction. The gray dashed line represents an arbitrary line of sight. Due to projection effects, the observed velocity of each spherical shell ($v_{\mathrm{obs}}$) corresponds to the velocity component parallel to the line of sight, and will be smaller than its intrinsic velocity (except for an impact parameter of zero, in which case they are the same). Thus, $v_{\mathrm{obs}}$ for a given spherical shell depends on the impact parameter of the line of sight, and the angle $\theta$ between the intersection of the spherical shell with the line of sight, and the $z$-axis (i.e., a higher $\theta$ implies that a lower fraction of the intrinsic radial velocity is parallel to the line of sight). The isovelocity surfaces are defined as the set of points within the outflow for which the observed velocity is identical, due to projection effects. Different values of $\gamma$ lead to different shapes of the isovelocity curves. In particular, $\gamma=1$ leads to perfectly vertical surfaces, meaning that the increase in radius (and thus, velocity) by moving upwards in the $y$ direction is exactly compensated by the decrease in $\cos(\theta)$, that is, a smaller fraction of the intrinsic velocity of the spherical shell is parallel to the line of sight, causing the projected velocity to remain unchanged. The most relevant aspect is that the isovelocity curves do not cross each other, implying that any line of sight intersects each isovelocity curve only once. Ultimately, the consequence of this is that for any arbitrary line of sight, different observed velocities map specific unique locations within the outflow (in the $z$ direction). This bijective relation between the velocity and physical space allows us to probe the outflow properties in a 3-D basis provided that we have sightlines for different locations in the $x-y$ plane (i.e., integral field data).
@@
$r^{2}-R_{0}^{2}\,\biggl{(}\frac{v_{\rm obs}}{v_{0}}\biggr{)}^{2/\gamma}\biggl{(}1-\biggl{[}\frac{\epsilon}{r}\biggr{]}^{2}\biggr{)}^{\frac{\gamma-1}{\gamma}}-\epsilon^{2}=0$ (3)
Thus, the optical depth ($\tau$) can be evaluated for each wavelength on the corresponding interaction surface, as a function of the physical properties of the wind and atomic constants, following Castor (1970):
$\tau(r)=\frac{\pi e^{2}}{mc}\,f_{lu}\,\lambda_{lu}\,n_{l}(r)\,\biggl{[}1-\frac{n_{u}\,g_{l}}{n_{l}\,g_{u}}\biggr{]}\frac{r/v}{1+\sigma\cos^{2}\phi}$ (4)
where $\phi$ is the angle between the velocity and the trajectory of the photon, $f_{ul}$ and $\lambda_{ul}$ correspond to the oscillator strength and wavelength for the $ul$ transition, respectively, and $\sigma=\frac{d\ln(v)}{d\ln(r)}-1$. Assuming the velocity law in Eq. 1 and neglecting stimulated emission (i.e., $\biggl{[}1-\frac{n_{u}\,g_{l}}{n_{l}\,g_{u}}\biggr{]}=1$), we get that the optical depth can be expressed as:
$\tau(r)=\frac{\pi e^{2}}{mc}\,f_{lu}\,\lambda_{lu}\,n_{0}\biggl{(}\frac{R_{0}}{r}\biggr{)}^{\gamma+2}\frac{r/v}{1+(\gamma-1)\cos^{2}\phi}$ (5)
Finally, defining:
$\tau_{0}=\frac{\pi e^{2}}{mc}\,f_{lu}\,\lambda_{lu}\,n_{0}\,\frac{R_{0}}{v_{0}}$ (6)
we can write the optical depth as:
$\tau(r)=\frac{\tau_{0}}{1+(\gamma-1)\,\cos^{2}\phi}\biggl{(}\frac{R_{0}}{r}\biggr{)}^{2\gamma+1}$ (7)

### v3

@@ 3.2 Integrated spectrum
In Fig. 2, we present a spectrum of the galaxy extracted by summing the data cube over a circular aperture of 1″ radius. The spectrum displays a blue stellar continuum with prominent Balmer absorption and strong nebular emission of [O II], [Ne III], the Balmer series, [O III], and other well-known lines. The spectrum shows Mg II to be mainly in absorption, but with significant redshifted P-Cygni emission humps as revealed by the inset panel in Fig. 2. Furthermore, Mg I $\lambda$2852 is detected in absorption.
@@
The emission line spectrum of #884 is that of a normal star-forming galaxy of moderate excitation level. The line widths also reveal no evidence for an AGN, nor do we find any AGN-characteristic emission line such as [Ne V] $\lambda$3426 in the spectrum. This is in agreement with the marginal detection ($5\sigma$, in the soft band only) as an X-ray source in the 7 Ms exposure of the Chandra observatory (Luo et al., 2017). The X-ray luminosity of $2\times 10^{41}$ erg s-1 is an order of magnitude below the limit usually adopted as the criterion for AGN, in line with the soft X-ray spectrum. Given the observed star formation rate, it is also consistent with the expected X-ray emission arising from stellar processes alone (Mineo et al., 2014). While these diagnostics do not exclude the possibility of a weak AGN in the center of the galaxy, all indicators suggest that any such source would have to be energetically subdominant.
Figure 3: Pseudo-narrowband Mg II image extracted from the MUSE data cube by summing over the spectral range of the doublet and filtering with a Gaussian of 1″ FWHM. Underlying in blue scale is the HST F775W image, with other galaxies labeled as in Fig. 1. The black solid contours are spaced logarithmically at 0.25 dex interval, starting at $\log_{10}(s)$ of $-18.5$, where $s$ is surface brightness in erg s-1 cm-2 arcsec-2. The lowest contour corresponds to a S/N of $\sim$1.5 in the filtered image. The dotted contours indicate negative measured values of $s$. The central hole in the Mg II surface brightness distribution is a result of the continuum subtraction, since #884 is a net absorber in its central spectrum. Similarly, the apparently negative region at the background galaxy BG1 is caused by the strong Mg II absorption in the continuum of BG1. The colored circumferences indicate the radii of the annular apertures used to extract radially resolved spectra in Sect. 4.2.
@@ 4.1 Mg II narrowband image
Similarly to the central depression, the Mg II emission in Fig. 3 also seems to be avoiding the galaxy BG1, and possibly also FG1. While for the former the explanation is again strong Mg II absorption in the underlying continuum of BG1, this is not possible for FG1 as the object is located in the foreground. However, any intrinsic stellar absorption feature in the FG1 spectrum that happens to be located within the Mg II narrowband range will have a similar effect. Since BG1 is sufficiently far away from #884 in terms of projected distance ($\sim 30$ kpc), we keep BG1 and its immediate surroundings outside the region where we model the Mg II emission (see Sect. 7).
At the upper edge of the image we see the bright foreground galaxy FG2 and the physical neighbor NB1. Most of the apparent signal in these spaxels is due to continuum-subtraction residuals from FG2 that are not restricted to the Mg II wavelength range. Similarly to BG1, this galaxy is also beyond the modeled region, to avoid our results being influenced by contaminant emission.
On the other hand, we find that NB1 is not surrounded by any large-scale extended Mg II emission (its apparent halo in Fig. 3 is an artifact generated by the residual of the continuum subtraction of FG2). The spectrum of NB1 shows only weak Mg II emission, with a 2-to-1 ratio between the two Mg II emission lines, suggesting a nebular origin. However, it is certainly possible that #884 and NB1 are in gravitational interaction, and that this interaction could influence the morphology of the Mg II halo through tidal forces, in the sense that it shows its largest extent along the direction toward NB1 (see, e.g., Leclercq et al., 2022).
Figure 4: Spatially resolved Mg II spectra of galaxy #884, extracted as the sum of spaxels in circular concentric ring-like apertures (circular aperture in the case of the central region). Each panel corresponds to a different annulus, as indicated in the label, except for the top-left panel, which corresponds to the spectrum of the full modeled FoV. In the frame of reference of the central galaxy, the apertures represent physical sizes of $<3$ kpc, $3-6$ kpc, $6-9$ kpc, $9-15$ kpc, and $15-26$ kpc. The error bars represent the ±$\pm 1\sigma$ uncertainty of the spectra. The nearly horizontal dotted lines indicate the approximate continuum level. The vertical thin dashed lines mark the observed wavelengths of the Mg II doublet at the systemic redshift of the galaxy.
@@ 4.2 Radially resolved spectra
In Fig. 4, we show spectra of the Mg II doublet, extracted in five disjoint concentric circular annular apertures, defined by the radii $<0\aas@@fstack{\prime\prime}4$, $0\aas@@fstack{\prime\prime}4-0\aas@@fstack{\prime\prime}8$, $0\aas@@fstack{\prime\prime}8-1\aas@@fstack{\prime\prime}2$, and $1\aas@@fstack{\prime\prime}2-2\aas@@fstack{\prime\prime}0$, or $<3$ kpc, $3-6$ kpc, $6-9$ kpc, $9-15$ kpc, and $15-26$ kpc, in the frame of reference of the central galaxy, as well as for the full modeled FoV. These apertures are shown in Fig. 3, following the same color scheme as in Fig. 4. The central spectrum (magenta) is dominated by absorption with only very weak P-Cygni emission humps. The emission then gets progressively stronger compared to the absorption when moving outward. At $1\aas@@fstack{\prime\prime}2<r<2\aas@@fstack{\prime\prime}0$ and beyond the absorption is essentially undetected.
@@ 5 Model
The main goal of this paper is to design a semianalytic model, with only a few parameters to avoid overfitting, that can reproduce key features observed of the Mg II halo of #884, in terms of the spatial distribution and line profiles of the Mg II emission and absorption, and that can be fitted to spatially resolved IFU observations. In the following subsections, we describe in detail the different components and other relevant aspects considered for our model, in order to achieve this goal. Figure 5 provides a flow chart that summarizes our modeling approach. Further in the paper, we describe the fit of our model to the data, aiming to unveil some of the key physics at play in the circumgalactic gas.
Figure 5: Flow chart summarizing our modeling and fitting scheme. Gray rounded boxes correspond to inputs for the model, blue rectangles correspond to intermediate outputs, and purple hexagons correspond to processes. The final output is indicated with a red ellipse. For a given set of parameters, a model cube is constructed on a spaxel-by-spaxel basis. The iteration over different individual spaxels is indicated with the yellow area. We used an MCMC routine to find the best-fitting parameters that match our data cube. The complete MCMC iteration over the parameter space is enclosed within the green (including the yellow) area. The inputs and intermediate outputs surrounded by a thick box are static, that is, they do not change over any of the iterations performed. In the top-left part of some boxes, we indicate the section where the concept inside the box is described. See Secs. 5 and 6 for a comprehensive description of our modeling and fitting approach.
@@ 5.1 Basic model
In this section, we describe the outflow model introduced by Scuderi et al. (1992) in the context of stellar winds and further developed by Scarlata & Panagia (2015), Carr et al. (2018), and Carr et al. (2021) to model the scattered resonant emission in expanding envelopes around galaxies. We also explain how we adapt this modeling scheme to integral field spectroscopic data. The basic model consists of a spherically symmetric source of radius $R_{0}$ isotropically emitting continuum radiation, surrounded by a spherical envelope of gas (wind) expanding from radius $R_{0}$ up to a terminal radius $R_{\rm max}$. The velocity at which the spherical envelope expands ($v$) increases with radius ($r$), according to a power law of exponent $\gamma$ such that
@@
where $v_{0}$ corresponds to the initial velocity of the wind (i.e., at $R_{0}$), and $v_{\mathrm{max}}$ is the terminal velocity of the wind at $R_{\mathrm{max}}$. A wind velocity that increases with radius is consistent with results from simulations (see, e.g., Cooper et al., 2008), and it is also consistent with the down-the-barrel observations reported by Martin & Bouché (2009), although different, more complex velocity profiles are possible (see, e.g., Murray et al., 2011). The density profile corresponding to Eq. 1, assuming mass conservation through the outflow, is
$n(r)=n_{0}\,\biggl{(}\frac{R_{0}}{r}\biggr{)}^{\gamma+2},$ (2)
@@
where $n_{0}$ is the gas density at $r=R_{0}$. Provided the velocity gradient in Eq. 1 is large, the photons produced by the central source will interact with the outflowing material only at the specific radius where the absorbing ions are at resonance (due to their Doppler shift), an assumption commonly known as the ”Sobolev” approximation. However, due to projection effects, the line-of-sight velocity $v_{\rm obs}$ measured by an observer for a given spherical shell at radius $r$ will differ from its real velocity, and will depend on the projected distance ($\epsilon$) of the line of sight to the central source (a.k.a., impact parameter), going from $v(r)$ at the center ($\epsilon=0$), to zero at $\epsilon=r$ (see also Scarlata & Panagia, 2015; Carr et al., 2018, for a more detailed explanation). On the other hand, it can be shown that the surfaces of constant $v_{\rm obs}$ do not cross each other, and hence, for any given line of sight, there is a bijective relation between frequency (or velocity) space and spatial location within the expanding envelope (because any sight line intersects different isovelocity surfaces only once). Figure 6 shows the shape of the (observed) isovelocity curves for $\gamma$ = 0.5, 1.0, and 1.5.
@@
Figure 6: Isovelocity curves in the spherical outflow model. The curves were calculated following Eq. 3 for three different values of $\gamma$, setting $R_{0}$ and $v_{0}$ to 1. The observer sits far away from the object toward the positive z-direction. The gray dashed line represents an arbitrary line of sight. Due to projection effects, the observed velocity of each spherical shell ($v_{\mathrm{obs}}$) corresponds to the velocity component parallel to the line of sight, and will be smaller than its intrinsic velocity (except for an impact parameter of zero, in which case they are the same). Thus, $v_{\mathrm{obs}}$ for a given spherical shell depends on the impact parameter of the line of sight, and the angle $\theta$ between the intersection of the spherical shell with the line of sight, and the $z$-axis (i.e., a higher $\theta$ implies that a lower fraction of the intrinsic radial velocity is parallel to the line of sight). The isovelocity surfaces are defined as the set of points within the outflow for which the observed velocity is identical, due to projection effects. Different values of $\gamma$ lead to different shapes of the isovelocity curves. In particular, $\gamma=1$ leads to perfectly vertical surfaces, meaning that the increase in radius (and thus, velocity) by moving upward in the $y$ direction is exactly compensated by the decrease in $\cos(\theta)$, that is, a smaller fraction of the intrinsic velocity of the spherical shell is parallel to the line of sight, causing the projected velocity to remain unchanged. The most relevant aspect is that the isovelocity curves do not cross each other, implying that any line of sight intersects each isovelocity curve only once. Ultimately, the consequence of this is that for any arbitrary line of sight, different observed velocities map specific unique locations within the outflow (in the $z$ direction). This bijective relation between the velocity and physical space allows us to probe the outflow properties on a 3-D basis provided that we have sight lines for different locations in the $x-y$ plane (i.e., integral field data).
@@
$r^{2}-R_{0}^{2}\,\biggl{(}\frac{v_{\rm obs}}{v_{0}}\biggr{)}^{2/\gamma}\biggl{(}1-\biggl{[}\frac{\epsilon}{r}\biggr{]}^{2}\biggr{)}^{\frac{\gamma-1}{\gamma}}-\epsilon^{2}=0.$ (3)
Thus, the optical depth ($\tau$) can be evaluated for each wavelength on the corresponding interaction surface as a function of the physical properties of the wind and atomic constants following Castor (1970):
$\tau(r)=\frac{\pi e^{2}}{mc}\,f_{lu}\,\lambda_{lu}\,n_{l}(r)\,\biggl{[}1-\frac{n_{u}\,g_{l}}{n_{l}\,g_{u}}\biggr{]}\frac{r/v}{1+\sigma\cos^{2}\phi},$ (4)
where $\phi$ is the angle between the velocity and the trajectory of the photon, $f_{ul}$ and $\lambda_{ul}$ correspond to the oscillator strength and wavelength for the $ul$ transition, respectively, and $\sigma=\frac{d\ln(v)}{d\ln(r)}-1$. Assuming the velocity law in Eq. 1 and neglecting stimulated emission (i.e., $\biggl{[}1-\frac{n_{u}\,g_{l}}{n_{l}\,g_{u}}\biggr{]}=1$), we get that the optical depth can be expressed as
$\tau(r)=\frac{\pi e^{2}}{mc}\,f_{lu}\,\lambda_{lu}\,n_{0}\biggl{(}\frac{R_{0}}{r}\biggr{)}^{\gamma+2}\frac{r/v}{1+(\gamma-1)\cos^{2}\phi}.$ (5)
Finally, defining
$\tau_{0}=\frac{\pi e^{2}}{mc}\,f_{lu}\,\lambda_{lu}\,n_{0}\,\frac{R_{0}}{v_{0}},$ (6)
we can write the optical depth as
$\tau(r)=\frac{\tau_{0}}{1+(\gamma-1)\,\cos^{2}\phi}\biggl{(}\frac{R_{0}}{r}\biggr{)}^{2\gamma+1}.$ (7)